OpenStudy (anonymous):

Let f(x) = x – 6 and g(x) = x^2 – x + 4. Find g(f(x)).

hartnn (hartnn):

to get g(f(x)) just replace 'x' in g(x) by f(x) = x-6 what u get ?

OpenStudy (hba):

g(f(x)=g(x-6)= Jusy plug in x-6 in place of x in x^2 – x + 4.

OpenStudy (anonymous):

(x-6)^2-(x-6)+4 and I'd solve this to get my answer?

hartnn (hartnn):

yep! just simplify it....

OpenStudy (hba):

Yes.

OpenStudy (anonymous):

k thx

OpenStudy (hba):

You are intelligento boy :)

hartnn (hartnn):

is it ?

OpenStudy (anonymous):

it's x^2-13x+46 and thx for helping :)

OpenStudy (hba):

Welcome :)

hartnn (hartnn):

oh, yes....you are intelligent!

OpenStudy (uri):

Give a medal to hartnn,Oh nvm i will give 3:

hartnn (hartnn):

thanks uri! but i am not here for medals.... he understood, thats my medal :)

OpenStudy (uri):

OMG THATS WHAT I WANTED TO HERE! EXACTLY THAT BRO! UREBA IS PROUD OF YOU! XDD